OpenStudy (anonymous):

Solve for x: square route of 3x^2-8 (end square route) + x=0

OpenStudy (anonymous):

To solve when there is a square root involved. you need to get the square root by itself on one side of the equation... So you would subtract x

OpenStudy (anonymous):

sqrt(3x^2 - 8) = -x

OpenStudy (anonymous):

Then you square both sides.

OpenStudy (anonymous):

sqrt(3x^2 - 8)^2 = (-x)^2 squaring a square root cancels each other out leaving you with what is left under the square root sign 3x^2 - 8 = x^2

OpenStudy (anonymous):

Make the equation = to 0 2x^2 - 8 = 0

OpenStudy (anonymous):

Factor 2(x^2 - 4) = 0 2(x + 2)(x - 2) = 0 x = -2 or x = 2

OpenStudy (anonymous):

Thanks!